2.Find the measure of the supplement of an angle that measures 72°
Found 3 solutions by rfer, user_dude2008, mananth:
Answer by rfer(16322)

180-72=108 degrees

Answer by user_dude2008(1862)

x+72=180
x+72-72=180-72
x=108 <--------------- Answer

Answer by mananth(16949)

sum of supplementary angles = 180 deg
one angle = 72
..
other angle = 180 -72
=108 deg